OpenStudy (anonymous):

If f(x) = 5x2 − 4 and g(x) = x − 2, what is (f o g)(-3)?

OpenStudy (anonymous):

\[f(x)=5x^{2}-4\] \[g(x)=x-2\] \[(fog)(x)=5(x-2)^{2}-4\] \[5(x ^{2}-4x+4)-4\]

OpenStudy (anonymous):

(fog)(x)=5x2-20x+20-4 (fog)(x)=5x2-20x+16 (fog)(-3)=5(-3)2-20(-3)+16 (fog)(-3)=45+60+16 (fog)(-3)=121

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my pleasure